how do i slove this system
3x+5y=-3
10y-2z=2
x=-z

Answer by Alan3354(69443)

how do i slove this system
3x+5y=-3
10y-2z=2
x=-z
------------
Sub -x for z in the 2nd eqn, then it's a 2 eqn system.
3x+5y=-3 --> 6x+10y=-6
10y+2x=2
--------
6x+10y=-6
2x+10y=2
-------- Subtract
4x = -8
--------
Take it from there.